OpenStudy (anonymous):

Enter the quadratic equation in standard form.2x2 - 3 = -4x - 1

OpenStudy (shubhamsrg):

the standard form of quadratic eqn is ax^2 + bx +c =0 will this help ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

but i just dont know this at all..

OpenStudy (shubhamsrg):

hmmn,,alright,,i'll solve this first one : bring all terms to one side we have 2x2 -3 +4x + 1 =0 collect like terms we have 2x2 +(1-3) +4x =0 => 2x2 -2 +4x arrange in form of ax2 + bx +c =0 i.e. degree wise.. so what should be the final ans ?

OpenStudy (anonymous):

2x^2 + 4x - 2 = 0

OpenStudy (shubhamsrg):

bingo!! ^_^

OpenStudy (anonymous):

Thanks :)